# A Study on Asynchronous Vote-based Blockchains

### v1

@@ 4.4.1 Vote validation
Honest nodes only accept votes that were cast following the honest model. We define the adversarial model in order to validate votes. An adversary must exhibit at least one of the two behaviours:
@@
When receiving a vote $V_{j}^{(k)}$, a node first constructs $G^{(k)}_{j}$ using the hash pointers in the vote. It then filters out any double-voting votes from $G_{j}^{(k)}$. Afterwards, it derives $VV_{j}^{(k)}$ and $VC_{j}^{(k)}$ and checks if $V_{j}^{(k)}$ is a compliant vote.
Filter out double voting votes: Double-voting is detected by checking if two or more votes from the same node at the same height exist in $G^{(k)}_{j}$. If double-voting by any node $j^{\prime}$ is detected in $G^{(k)}_{j}$, only the votes prior to the first fork of the vote chain of node $j^{\prime}$ is counted in $VV_{j}^{(k)}$. Fig. 2 depicts a scenario of a four-node system. When a node receives vote $V^{(5)}_{4}$, it reconstructs $G^{(5)}_{4}$ as shown in the figure. From this graph, $V_{4}^{(5)}$ is linked to $V_{1}^{(4)}$, $V_{2}^{(4)}$ and its own $V_{4}^{(4)}$. Analysis of the paths in the graph reveals a double-voting for $V_{1}^{(2)}$, identified through distinct paths $V_{4}^{(5)}\rightarrow V_{1}^{(4)}\rightarrow V_{1}^{(3)}\rightarrow V_{1}^{(2)}$ and $V_{4}^{(5)}\rightarrow V_{2}^{(4)}\rightarrow V_{2}^{(3)}\rightarrow V_{1}^{(2)}$, etc. Any node receives $V^{(5)}_{4}$ can reconstruct an identical $G^{(5)}_{4}$. Following the detection of double-voting votes, all subsequent votes related to the anomaly are removed from $VV_{4}^{(5)}$ and will not be counted in $VC_{4}^{(5)}$.
@@
Verify the compliance of voting rules: If a vote is not classified as a compliant vote, it is classified as a non-compliant vote. $V_{j}^{(k)}$ is a compliant vote if it satisfies all the following points: (1) $V_{j}^{(k)}$ was cast for a block at consensus height $k$; (2) $G_{j}^{(k)}$ does not contain non-compliant votes; (3) $V_{j}^{(k)}$ was cast for a block $B_{k}$, which is in the branch rooted by $B^{\prime}_{k^{\prime}}$ of consensus height $k^{\prime}\in[1,k-1]$. $B^{\prime}_{k^{\prime}=1}$ originated from the genesis block. Each other block $B^{\prime\prime}_{k^{\prime}}$ of consensus height $k^{\prime}$, which is in the same branch as $B^{\prime}_{k^{\prime}}$ (they both stem from $B^{\prime}_{k^{\prime}-1}$ of consensus height $k^{\prime}-1$), fulfills
$VC_{j}^{(k)}[B^{\prime}_{k^{\prime}}]\geq VC_{j}^{(k)}[B^{\prime\prime}_{k^{\prime}}]$ (1)
Alg. 2 shows the pseudocode for vote validation. Lem. 2 shows that all nodes must vote in the most supported branch suggested by the vote counts.
@@
4:    $VV_{j}^{(k)}\leftarrow$ DeriveVoteView($G_{j}^{(k)}$)
5:    $VC_{j}^{(k)}\leftarrow$ DeriveVoteCounte($VV_{j}^{(k)}$)
6:    Verify if $V^{(k)}_{j}$ follows voting rules using $VC_{j}^{(k)}$ and $G_{j}^{(k)}$
@@
Assume that the lemma is false. That is, assume there exists an adversarial node $j$ whose vote $V_{j}^{(k)}$ for a block stemming from $B_{k^{\prime}}$ at height $k$ is linked by an honest node at height $k+1$, even though $B_{k^{\prime}}$ is not stem from the most supported branch according to $VC_{j}^{(k)}$.
@@
Under the voting rules, a compliant vote is only linked to a vote if it satisfies the criteria outlined for a compliant vote. According to the criteria, a node $i$ will only consider $V_{j}^{(k)}$ if: - The block $B_{k^{\prime}}$ voted on by $j$ has the highest vote count among all blocks at height $k^{\prime}$ of the same branch and this branch stems from the most supported branch, according to $VC_{j}^{(k)}$.
@@
  • $V_{j}^{(k)}$ adhered to the honest voting criteria 2, contradicting the assumption that $B_{k^{\prime}}$ was not the root of the most supported branch stemming from the most supported branch.
  • Therefore, if $B_{k^{\prime}}$ was indeed not the root of the most supported branch stemming from the most supported branch based on $VC_{j}^{(k)}$, the honest nodes would have disregarded $V_{j}^{(k)}$ due to the enforcement of compliance checks $VC_{j}^{(k)}[B_{k^{\prime}}]<VC_{j}^{(k)}[B^{\prime\prime}_{k^{\prime}}]$ for any other block $B^{\prime\prime}_{k^{\prime}}$ with higher support.
@@
Conclusion: Given the system’s design to filter non-compliant votes and enforce the honest model, it is contradictory to assume a scenario where an adversarial vote for a less supported branch gets linked by an honest node. So, by contradiction, the lemma holds under the protocol’s operational assumptions and security model. Thus, an adversary must vote for the most supported branch to their local knowledge to have their vote considered by others.
@@ 4.5 Reaching Consensus
$\forall j\in NG,\quad\quad\quad VC_{j}^{(k)}[B_{k^{\prime}}]-t>$ (2)
$VC_{j}^{(k)}[B^{\prime}_{k^{\prime}}(j)]+VC_{j}^{(k)}[U_{k}]-VC_{j}^{(k)}[U_{k^{\prime}}]-1$
@@
  • $B^{\prime}_{k^{\prime}}(j)$ is the second most supported block of the consensus height $k^{\prime}$ according to $VC_{j}^{(k)}$. If there is no block other than $B_{k^{\prime}}$, then $VC_{j}^{(k)}[B^{\prime}_{k^{\prime}}(j)]=0$.
  • $VC_{j}^{(k)}[U_{k^{\prime}}]$ is the number of votes from height $1$ to $k^{\prime}-1$ (inclusive) that are yet to be received by node $j$ according to $G_{j}^{(k)}$  [^6] There are $N\times(k^{\prime}-1)$ votes from height $1$ to $k^{\prime}-1$, one can locate all the received votes from $G_{j}^{(k)}$, then calculate the missing vote number.
@@
9:             $voteForB_{k^{\prime}}\leftarrow$ VoteCount($B_{k^{\prime}}$, $VV_{j}^{(k)}$)
10:             $voteForB^{\prime}_{k^{\prime}}\leftarrow$ VoteCount($B^{\prime}_{k^{\prime}}(j)$, $VV_{j}^{(k)}$)
@@
34:function VoteCount($B_{k^{\prime}}$, $VV_{j}^{(k)}$)
35:    return Count of votes for $B_{k^{\prime}}$ and blocks stemmed from $B_{k^{\prime}}$ in $VV_{j}^{(k)}$
@@
Rationale behind Eq. 2: Assume that a block $B_{k^{\prime}}$ of consensus height $k^{\prime}$ is consensually accepted at height $k>k^{\prime}$ (Eq. 2 is satisfied). When $VC_{j}^{(k)}[B_{k^{\prime}}]>VC_{j}^{(k)}[B^{\prime}_{k^{\prime}}(j)]+VC_{i}^{(k)}[U_{k}]-VC_{i}^{(k)}[U_{k^{\prime}}]$, any node $j$ in $NG$ would have voted for the branch of $B_{k^{\prime}}$ at height $k$. Failure to do so would imply adversarial characteristic (C2) and is therefore not in $NG$. Despite this, network delays might cause some node $i$ in $NG$ to send votes for height $k+1$ before receiving some other votes at $k$ from nodes in $NG$. This may potentially affect node $i$’s vote for height $k+1$ because in $G_{i}^{(k+1)}$ there could be less than $N-f$ nodes voted for blocks stemming from $B_{k^{\prime}}$ at $k$. To ensure continued support for $B_{k^{\prime}}$ at $k+1$, it must be demonstrated that for any vote of height $k+1$ linked to any $N-f$ votes at vote height $k$, $B_{k^{\prime}}$ remains the root of the most supported branch of $k^{\prime}$ in its vote count and this support is greater than the second most supported plus the missing votes.
In the worst-case scenario, only $f+1$ votes of height $k$ from nodes in $NG$ are linked by a new vote of node $i$ of height $k+1$. This is because $N=3f+1$ and $NG$ contains $12f+$ 1 nodes. As mentioned, nodes in $NG$ must have voted in $B_{k^{\prime}}$’s branch at height $k$. After receiving $N-f$ votes at height $k$, if $VC_{i}^{(k)}[B_{k^{\prime}}]+f+1>VC_{i}^{(k)}[U_{k}]-VC_{i}^{(k)}[U_{k^{\prime}}]+VC_{i}^{(k)}[B^{\prime}_{k^{\prime}}(i)]+f$, node $i$ will still vote in $B_{k^{\prime}}$’s branch at $k+1$. This is because, if node $i$ discovers all missing votes and finds that they support $B^{\prime}_{k^{\prime}}(i)$, and there are $f$ votes supporting $B^{\prime}_{k^{\prime}}(i)$ in $G_{i}^{(k+1)}$, the branch rooted by block $B_{k^{\prime}}$ remains the most supported.
If adversarial characteristic (C1) is detected in the future, double votes are not counted. Assuming the votes were counted in $VC_{i}^{(k)}[B_{k^{\prime}}]$ already, then it should satisfy $VC_{i}^{(k)}[B^{\prime}_{k^{\prime}}]+f+1-t>VC_{i}^{(k)}[U_{k}]-VC_{i}^{(k)}[U_{k^{\prime}}]+VC_{i}^{(k)}[B^{\prime}_{k^{\prime}}(i)]+f$, where $t=(k-k^{\prime})\times f$.
Thus, once Eq. 2 is satisfied, nodes cannot vote for blocks in other branches, establishing the uniqueness of acceptance for block $B_{k^{\prime}}$. Lem. 4.5 is a formal statement of the safety of this algorithm.
@@
    $VC_{j}^{(k)}[B_{k^{\prime}}]-t>VC_{j}^{(k)}[B^{\prime}_{k^{\prime}}(j)]+VC_{j}^{(k)}[U_{k}]-VC_{j}^{(k)}[U_{k^{\prime}}]-1$
@@
4.6 Livenss of Consensus Derivation
To guarantee consensus-derivation liveness, we must exhibit that the number of blocks ($m$) at a specific consensus height supported by nodes—by casting votes for blocks stemming from them—monotonically decreases until all the nodes have voted for the same branch. The asynchronous network stipulates that the honest message delay is arbitrary but finite. It does not address to what extent adversaries can manipulate these delays. Here, we detail different situations.
@@ 4.6.1 Practical liveness
It is generally known that asynchronous BFT consensus protocols require unpredictable randomness so that an adversary can only adapt to the randomness when it is too late to influence the outcome.
Random message delays introduce entropy into the system. In our design, each node derives its voting decision from the first $N-f$ compliant votes of the previous height it has received, akin to utilizing a weak common coin [1], assuming that an adversary has no control over which $N-f$ votes each node will receive first. This consequently reduces $m$ when some branches no longer remain the most supported in any vote’s vote count. Ultimately, only one branch remains. This approach possesses several key properties:
@@
Lemma (Normal scenario) The amortized rounds of voting before convergence is within $\log_{2}(m)$ rounds.
In a scenario where no value has an overwhelming majority initially, the convergence to a single value is driven by the randomness of the sampled majorities. The expected rounds to converge, assuming very unfavorable conditions (e.g., a close initial split of supports and a strong adaptive adversary observing others’ votes before casting their own at each round), is $\log_{2}(m)$ rounds as the process of voting and random sampling amplifies any emerging majority.
@@
Appx. B shows a simple simulation that faithfully represents the voting rounds before convergence. With $N=100$ and $m=100$, convergence happens within $4$ voting rounds.

### v2

@@ 3.4.1 Vote validation
Honest nodes only accept votes that were cast following the honest model. We define the adversarial model in order to validate votes. A Byzantine node must exhibit at least one of the two behaviours:
@@
When receiving a vote $V_{j}^{(k)}$, a node first constructs $G^{(k)}_{j}$ using the hash pointers in the vote. It then filters out any double-voting votes from $G_{j}^{(k)}$. Afterwards, it derives $\operatorname{VV}_{j}^{(k)}$ and $\operatorname{VC}_{j}^{(k)}$ and checks if $V_{j}^{(k)}$ is a compliant vote.
Filter out double voting votes: Double-voting is detected by checking if two or more votes from the same node at the same height exist in $G^{(k)}_{j}$. If double-voting by any node $j^{\prime}$ is detected in $G^{(k)}_{j}$, only the votes prior to the first fork of the vote chain of node $j^{\prime}$ is counted in $\operatorname{VV}_{j}^{(k)}$. Fig. 2 depicts a scenario of a four-node system. When a node receives vote $V^{(5)}_{4}$, it reconstructs $G^{(5)}_{4}$ as shown in the figure. From this graph, $V_{4}^{(5)}$ is linked to $V_{1}^{(4)}$, $V_{2}^{(4)}$ and its own $V_{4}^{(4)}$. Analysis of the paths in the graph reveals a double-voting for $V_{1}^{(2)}$, identified through distinct paths $V_{4}^{(5)}\rightarrow V_{1}^{(4)}\rightarrow V_{1}^{(3)}\rightarrow V_{1}^{(2)}$ and $V_{4}^{(5)}\rightarrow V_{2}^{(4)}\rightarrow V_{2}^{(3)}\rightarrow V_{1}^{(2)}$, etc. Any node receives $V^{(5)}_{4}$ can reconstruct an identical $G^{(5)}_{4}$. Following the detection of double-voting votes, all subsequent votes related to the anomaly are removed from $\operatorname{VV}_{4}^{(5)}$ and will not be counted in $\operatorname{VC}_{4}^{(5)}$.
@@
Verify the compliance of voting rules: If a vote is not classified as a compliant vote, it is classified as a non-compliant vote. $V_{j}^{(k)}$ is a compliant vote if it satisfies all the following points: (1) $V_{j}^{(k)}$ was cast for a block at consensus height $k$; (2) $G_{j}^{(k)}$ does not contain non-compliant votes; (3) $V_{j}^{(k)}$ was cast for a block $B_{k}$, which is in the branch rooted by $B_{k^{\prime}}$ of consensus height $k^{\prime}\in[1,k-1]$. $B_{k^{\prime}=1}$ originated from the genesis block. Each other block $B^{\prime}_{k^{\prime}}$ of consensus height $k^{\prime}$, which is in the same branch as $B_{k^{\prime}}$ (they both stem from $B_{k^{\prime}-1}$ of consensus height $k^{\prime}-1$), fulfills
$\operatorname{VC}_{j}^{(k)}[B_{k^{\prime}}]\geq\operatorname{VC}_{j}^{(k)}[B^{\prime}_{k^{\prime}}]$ (1)
Alg. 2 shows the pseudocode for vote validation. Lem. 2 shows that all nodes must vote in the most supported branch suggested by the Vote Count of their votes. Further justifications for compliance checking see Appx. C.3.
@@
4:    $\operatorname{VV}_{j}^{(k)}\leftarrow$ DeriveVoteView($G_{j}^{(k)}$)
5:    $\operatorname{VC}_{j}^{(k)}\leftarrow$ DeriveVoteCount($\operatorname{VV}_{j}^{(k)}$)
6:    Verify if $V^{(k)}_{j}$ follows voting rules using data in $\operatorname{VC}_{j}^{(k)}$ and $G_{j}^{(k)}$
@@
Assume that the lemma is false. That is, assume there exists an adversarial node $j$ whose vote $V_{j}^{(k)}$ for a block stemming from $B_{k^{\prime}}$ at height $k$ is linked by an honest node at height $k+1$, even though $B_{k^{\prime}}$ is not stem from the most supported branch according to $\operatorname{VC}_{j}^{(k)}$.
@@
Under the voting rules, a compliant vote is only linked to a vote if it satisfies the criteria outlined for a compliant vote. According to the criteria, a node $i$ will only consider $V_{j}^{(k)}$ if: - The block $B_{k^{\prime}}$ voted on by $j$ has the highest vote count among all blocks at height $k^{\prime}$ of the same branch and this branch stems from the most supported branch, according to $\operatorname{VC}_{j}^{(k)}$.
@@
  • $V_{j}^{(k)}$ adhered to the honest voting criteria 2, contradicting the assumption that $B_{k^{\prime}}$ was not the root of the most supported branch stemming from the most supported branch according to $\operatorname{VC}_{j}^{(k)}$
  • Therefore, if $B_{k^{\prime}}$ was indeed not the root of the most supported branch stemming from the most supported branch based on $\operatorname{VC}_{j}^{(k)}$, the honest nodes would have disregarded $V_{j}^{(k)}$ due to the enforcement of compliance checks $\operatorname{VC}_{j}^{(k)}[B_{k^{\prime}}]<\operatorname{VC}_{j}^{(k)}[B^{\prime}_{k^{\prime}}]$ for any other block $B^{\prime}_{k^{\prime}}$ with higher support.
@@
Conclusion: Given the system’s design to filter non-compliant votes and enforce the honest model, it is contradictory to assume a scenario where an adversarial vote for a less supported branch gets linked by an honest node. So, by contradiction, the lemma holds under the protocol’s operational assumptions and security model. Thus, a Byzantine node must vote for the most supported branch to their local knowledge to have their vote considered by others.
@@ 3.5 Reaching Consensus
$\forall j\in NG:\quad\quad\quad\operatorname{VC}_{j}^{(k)}[B_{k^{\prime}}]-t>$ (2)
$\operatorname{VC}_{j}^{(k)}[B^{\prime}_{k^{\prime}}(j)]+\operatorname{VC}_{j}^{(k)}[U_{k}]-\operatorname{VC}_{j}^{(k)}[U_{k^{\prime}}]-1$
@@
  • $B^{\prime}_{k^{\prime}}(j)$ is the second most supported block of the consensus height $k^{\prime}$ according to $\operatorname{VC}_{j}^{(k)}$. If there is no block other than $B_{k^{\prime}}$, then $\operatorname{VC}_{j}^{(k)}[B^{\prime}_{k^{\prime}}(j)]=0$.
  • $\operatorname{VC}_{j}^{(k)}[U_{k^{\prime}}]$ is the number of votes from height $1$ to $k^{\prime}-1$ (inclusive) that are yet to be received by node $j$ according to $G_{j}^{(k)}$  [^6] There are $N\times(k^{\prime}-1)$ votes from height $1$ to $k^{\prime}-1$, one can locate all the received votes from $G_{j}^{(k)}$, then calculate the missing vote number.
@@
9:             $voteForB_{k^{\prime}}\leftarrow$ VoteCount($B_{k^{\prime}}$, $\operatorname{VV}_{j}^{(k)}$)
10:             $voteForB^{\prime}_{k^{\prime}}\leftarrow$ VoteCount($B^{\prime}_{k^{\prime}}(j)$, $\operatorname{VV}_{j}^{(k)}$)
@@
34:function VoteCount($B_{k^{\prime}}$, $\operatorname{VV}_{j}^{(k)}$)
35:    return Count of votes for $B_{k^{\prime}}$ and blocks stemmed from $B_{k^{\prime}}$ in $\operatorname{VV}_{j}^{(k)}$
@@
Rationale behind Eq. 2: Assume that a block $B_{k^{\prime}}$ of consensus height $k^{\prime}$ is consensually accepted at height $k>k^{\prime}$ (Eq. 2 is satisfied). When $\operatorname{VC}_{j}^{(k)}[B_{k^{\prime}}]>\operatorname{VC}_{j}^{(k)}[B^{\prime}_{k^{\prime}}(j)]+\operatorname{VC}_{i}^{(k)}[U_{k}]-\operatorname{VC}_{i}^{(k)}[U_{k^{\prime}}]$, any node $j$ in $NG$ would have voted for the branch of $B_{k^{\prime}}$ at height $k$. Failure to do so would imply adversarial characteristic (C2) and is therefore not in $NG$. Despite this, network delays might cause some node $i$ in $NG$ to send votes for height $k+1$ before receiving some other votes at $k$ from nodes in $NG$. This may potentially affect node $i$’s vote for height $k+1$ because in $G_{i}^{(k+1)}$ there could be less than $N-f$ nodes voted for blocks stemming from $B_{k^{\prime}}$ at $k$. To ensure continued support for $B_{k^{\prime}}$ at $k+1$, it must be demonstrated that for any vote of height $k+1$ linked to any $N-f$ votes at vote height $k$, $B_{k^{\prime}}$ remains the root of the most supported branch of $k^{\prime}$ in its vote count and this support is greater than the second most supported plus the missing votes.
In the worst-case scenario, only $f+1$ votes of height $k$ from nodes in $NG$ are linked by a new vote of node $i$ of height $k+1$. This is because $N=3f+1$ and $NG$ contains $12f+$ 1 nodes. As mentioned, nodes in $NG$ must have voted in $B_{k^{\prime}}$’s branch at height $k$. After receiving $N-f$ votes at height $k$, if $\operatorname{VC}_{i}^{(k)}[B_{k^{\prime}}]+f+1>\operatorname{VC}_{i}^{(k)}[U_{k}]-\operatorname{VC}_{i}^{(k)}[U_{k^{\prime}}]+\operatorname{VC}_{i}^{(k)}[B^{\prime}_{k^{\prime}}(i)]+f$, node $i$ will still vote in $B_{k^{\prime}}$’s branch at $k+1$. This is because, if node $i$ discovers all missing votes and finds that they support $B^{\prime}_{k^{\prime}}(i)$, and there are $f$ votes supporting $B^{\prime}_{k^{\prime}}(i)$ in $G_{i}^{(k+1)}$, the branch rooted by block $B_{k^{\prime}}$ remains the most supported.
If adversarial characteristic (C1) is detected in the future, double votes are not counted. Assuming the votes were counted in $\operatorname{VC}_{i}^{(k)}[B_{k^{\prime}}]$ already, then it should satisfy $\operatorname{VC}_{i}^{(k)}[B_{k^{\prime}}]+f+1-t>\operatorname{VC}_{i}^{(k)}[U_{k}]-\operatorname{VC}_{i}^{(k)}[U_{k^{\prime}}]+\operatorname{VC}_{i}^{(k)}[B^{\prime}_{k^{\prime}}(i)]+f$, where $t=(k-k^{\prime})\times f$.
Thus, once Eq. 2 is satisfied, nodes cannot vote for blocks in other branches, establishing the uniqueness of acceptance for block $B_{k^{\prime}}$. Lem. 3.5 is a formal statement of the safety of this algorithm.
@@
    $\operatorname{VC}_{j}^{(k)}[B_{k^{\prime}}]-t>\operatorname{VC}_{j}^{(k)}[B^{\prime}_{k^{\prime}}(j)]+\operatorname{VC}_{j}^{(k)}[U_{k}]-\operatorname{VC}_{j}^{(k)}[U_{k^{\prime}}]-1$
@@
For better understanding, Appx. E shows an online visualization demonstration of our protocol.
3.6 Liveness of Consensus Derivation
To guarantee consensus-derivation liveness, we must exhibit that the number of blocks ($m$) at a specific consensus height supported by nodes—by casting votes for blocks stemming from them—monotonically decreases until all the nodes have voted for the same branch. The asynchronous network stipulates that the honest message delay is arbitrary but finite. It does not address to what extent Byzantine nodes can manipulate these delays. Here, we detail different situations.
@@ 3.6.1 Practical liveness
It is generally known that asynchronous BFT consensus protocols require unpredictable randomness so that the adversary can only adapt to the randomness when it is too late to influence the outcome.
Random message delays introduce entropy into the system. In our design, each node derives its voting decision from the first $N-f$ compliant votes of the previous height it has received, akin to utilizing a weak common coin [1], assuming that the adversary has no control over which $N-f$ votes each node will receive first. This consequently reduces $m$ when some branches no longer remain the most supported in any vote’s vote count. Ultimately, only one branch remains. This approach possesses several key properties:
@@
[end,restate]Lemma (Normal case) Consider $N$ nodes voting for one of $m\leq N$ distinct input values. In each round, a node can observe $N-f$ votes from the previous round, and each vote is recursively linked to the $N-f$ votes from the previous round. The process of convergence can be modeled as a discrete-time Markov chain (DTMC) with $m$ states, representing the number of remaining input values. In this model: Worst Case: The convergence time for all nodes to vote for the same input value is $O(m)$ rounds. Amortized Case: In the average case, the system converges in $O(\log m)$ rounds.
{proofEnd}
We model the process of convergence as a discrete-time Markov chain (DTMC) with state space $\{1,2,\dots,m\}$, where each state $x_{r}=x$ represents the number of distinct input values remaining at round $r$. The absorbing state is $x=1$, where consensus is reached, and no further transitions occur.
State Space and Transitions:
  • The process begins at state $x_{r}=m$, where all $m$ distinct input values are still in play.
  • The system transitions between states based on whether one or more input values are eliminated in each round. A transition from $x_{r}=x$ to $x_{r+1}=x-j$ represents the elimination of $j$ input values in that round, where $j\in\{1,2,\dots,x-1\}$.
  • The absorbing state is $x=1$, where all nodes vote for the same input value and convergence is reached.
Transition Probabilities: Let $P(x\to x)$ be the probability that no progress is made in a round, meaning no input value is eliminated. Let $P(x\to x-j)$ be the probability that $j$ input values are eliminated in a round.
  1. 1.
    No Progress Probability $P(x\to x)$: This occurs when the nodes nearly equal supporting the remaining $x$ input values, and no clear majority emerges. The probability of no progress is higher when the votes are more evenly distributed. Based on binomial distribution approximations, a very lost upper bound probability can be written as:
    $P(x\to x)\approx\frac{1}{\sqrt{2\pi\frac{N-f}{4}}}$
    The rational behind this probability is that consider the case where there are two equally supported branches rooted by block $A$ and $B$, and each has $\frac{N}{2}$ votes. Each node samples $N-f$ votes from this set. Afterwards, the probability that exactly $\frac{N}{2}$ nodes vote for each branch (thus no change in the poll) can be approximated using the binomial distribution:
    $p(o)=\binom{N-f}{o}\left(\frac{1}{2}\right)^{o}\left(\frac{1}{2}\right)^{(N-f-o)}=\binom{N-f}{o}\left(\frac{1}{2}\right)^{N-f}$
    where $o=\frac{N-f}{2}$. The probability (very loose upper bound) is thus approximately:$p_{(\text{no progress})}\approx\frac{1}{\sqrt{2\pi\frac{N-f}{4}}}$.
  2. 2.
    Progress Probability $P(x\to x-j)$: The probability that $j$ input values are eliminated in the round depends on how biased the vote distribution is. As $x$ decreases (fewer input values remain), the likelihood of eliminating multiple input values at once increases if the vote distribution skews towards one or more dominant input values.
    The probabilities for eliminating $j$ input values in a single round can be modeled as a decreasing geometric-like distribution where the probability of eliminating more input values declines as $j$ increases:
    $P(x\to x-j)=\alpha_{j}\cdot P(\text{progress}),\quad\sum_{j=1}^{x-1}\alpha_{j}=1$
    where $\alpha_{j}$ is a weight assigned to the probability of eliminating $j$ input values. For simplicity, these weights can reflect the natural tendency that eliminating fewer input values is more likely than eliminating more in each round.
Transition Matrix: The transition matrix $P$ for this Markov chain is defined as follows, where $P_{ij}$ represents the probability of transitioning from state $i$ to state $j$:
$\small P=\begin{pmatrix}1&0&0&\dots&0&0\\ P(2\to 1)&P(2\to 2)&0&\dots&0&0\\ P(3\to 1)&P(3\to 2)&P(3\to 3)&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ P(m\to 1)&P(m\to 2)&P(m\to 3)&\dots&P(m\to m-1)&P(m\to m)\\ \end{pmatrix}$
Expected Time to Absorption: The goal is to calculate the expected time to reach the absorbing state $x=1$, starting from state $x=m$. Let $t_{x}$ represent the expected number of rounds to reach state $x=1$ starting from state $x$. The recurrence relation for $t_{x}$ is:
$t_{x}=\frac{P(x\to x)+\sum_{j=1}^{x-1}P(x\to x-j)+\sum_{j=1}^{x-1}P(x\to x-j)\cdot t_{x-j}}{1-P(x\to x)}$
This equation indicates that the expected time $t_{x}$ depends on either staying in the same state (with probability $P(x\to x)$) or transitioning to lower states, reducing multiple input values at once. The summation term accounts for all possible transitions, where $j$ input values are eliminated in a round.
Rearranging the terms:
$t_{x}=\frac{1+\sum_{j=1}^{x-1}P(x\to x-j)\cdot t_{x-j}}{P(\text{progress})}$
By solving this recurrence relation starting from $t_{1}=0$ (the absorbing state), we can calculate the expected time to convergence.
Worst-Case Convergence Time: In the worst case, when the votes are nearly uniformly distributed, $P(x\to x)$ remains high, and progress is slow. The recurrence relation shows that when $P(x\to x)$ is significant, the expected time $t_{x}$ increases linearly with $x$, leading to an overall convergence time of $O(m)$.
Amortized Case Convergence Time: In the average case, as $x$ decreases, the vote distribution becomes less uniform, and the probability of eliminating multiple input values in a round increases. Once a majority input value emerges, recursive vote linking amplifies its support, leading to exponential convergence. The expected time to consensus in this case is $O(\log m)$.
@@
Appx. D shows a simple simulation that faithfully represents the voting rounds before convergence. With $N=100$ and $m=100$, convergence happens within $4$ voting rounds.